# Graph of AI Ideas: Leveraging Knowledge Graphs and LLMs for AI Research Idea Generation

### v1

@@ 4.1 Setup
Multi-turn Dialogue 0.3119
Structured Thinking (ours) 0.5271
Table 2: The impact of SFT dataset on evaluation model.
Figure 4: Evaluation results of idea generation.
Figure 5: Case study for the idea generation result of our GoAI compared with Chain of Ideas Li et al. (2024) in the same topic “Eliminate Hallucinations in Large Language Models”. The paper number from 1 to 6 represent Qi et al. (2021), Min et al. (2020), Zhou et al. (2021), Ahn et al. (2022), Izacard et al. (2022), and Chang et al. (2024), respectively.
4.2 Main Results
4.2.1 Human-Model Agreements of Evaluation
Table 2 presents the Pearson correlation coefficient between our evaluation model and human evaluations. The results show that our evaluation model maintains a good positive correlation with human reviewers’ assessments, making it a valuable tool for providing feedback in the idea generation process.
4.2.2 Novelty of Generated Ideas
Figure 4 and Table 1 display a comparison of the creative idea generation results between our proposed method and other idea generation approaches. Overall, compared to all other automated methods, our GoAI demonstrates superior performance across evaluations. Notably, GoAI excels in generating ideas with novelty and significance, highlighting its exceptional capability in creative idea generation. In other areas, considering that no targeted optimization was performed, the gap between GoAI and other methods is expected. Nevertheless, even in these aspects, GoAI still exhibits strong performance, allowing it to outperform other methods in terms of average scores.
4.3 Ablation Study for Evaluation Model
To investigate the impact of the structured thinking process SFT training dataset on the evaluation model, we compared it with models trained using raw data directly and using datasets in a multi-turn dialogue format. The evaluation metric used is the Pearson correlation coefficient with human ratings. The dataset trained with raw data takes the form of a QA question-answering format, where the input is the title and abstract of a paper, and the output is the novelty score of the paper. The multi-turn dialogue format dataset consists of several rounds of QA about the paper’s summary, strengths and weaknesses, and score. Table 2 presents the Pearson correlation coefficient results for the models trained with different datasets. The results show that using the structured CoT thinking process significantly improves the correlation between the evaluation model and human assessments, consistent with the findings in Xu et al. (2025).
4.4 Case Study
Figure 5 presents a case study comparing our proposed GoAI method with the Chain of Ideas Li et al. (2024). The topic used for generating ideas is “Eliminate Hallucinations in Large Language Models,” and the starting point for constructing path analysis of research trends is the same paper, “Retrieval Augmentation Reduces Hallucination in Conversation.” From the paper paths, it can be observed that our method, by analyzing the semantic information of citations, is able to retrieve papers more closely related to the topic. The generated idea also aligns more closely with the paper path, indicating that the idea generation comes from a thorough analysis of the development trends in the field. In contrast, the paths constructed by the Chain of Ideas method include a paper that is unrelated to the topic being studied, titled “cTBLS: Augmenting Large Language Models with Conversational Tables.” The generated idea is also unrelated to the papers in the path, indicating that the papers retrieved by this method are easily influenced by irrelevant citations, and the generated idea may originate from the model’s own pre-trained knowledge or hallucinations.
5 Conclusion
In this paper, we introduced the Graph of AI Ideas (GoAI), a framework designed to enhance the capability of LLMs to generate research ideas in the field of AI. By organizing ideas from existing papers into a structured graph, GoAI effectively captures the progressive relationships and incremental developments between studies. The use of an LLM-based agent for path searching in the graph helps LLMs to understand and analyze current research progress, uncover connections between existing studies, and thereby enhance their creative potential.
To evaluate and refine the automatically generated ideas, we also proposed GoAI-CoT-Reviewer, an evaluation model with structured thinking capabilities. This model evaluates the novelty of ideas through a reasoning process similar to that of human reviewers, providing feedback highly consistent with human assessments. Experimental results demonstrate that GoAI consistently outperforms other methods and is capable of generating ideas comparable to human creativity.
Ethical Statement
This study constructs a knowledge graph using publicly available research paper data to generate novel ideas. All data were sourced from open academic repositories, with no personal or sensitive information involved. Data processing followed the respective terms of use, and no privacy concerns arose. The authors declare no conflicts of interest, and all analyses were conducted transparently using automated methods. The study complies with ethical guidelines for responsible data use in academic research.

### v2

@@ Introduction
Multi-turn Dialogue 0.4251
Structured Thinking (ours) 0.6810
Table 2: Pearson correlation coefficients between different evaluation models and human assessment.
Idea Studio
We present Idea Studio as a demonstrative application of the GoAI framework. For an AI major student, Idea Studio provides structured prompts requiring them to conduct research and formulate novel ideas based on a specific topic—for instance, “Enhancing the Reasoning Ability of Large Language Models.” By leveraging GoAI to generate research trajectories and personalized learning paths, students gain access to the developmental trends and key literature within the selected domain, enabling them to independently acquire the necessary prerequisite knowledge. Upon completing the study of relevant concepts and reviewing the associated papers, students propose original ideas informed by the identified trends, which are then evaluated and iteratively refined through the GoAI-CoT-Reviewer. This process not only serves as an empirical evaluation of GoAI’s practical utility but also functions as a pedagogical tool to support the everyday learning and research activities of AI students.
Experiments
Setup
We use the Deepseek-V3 API as the backend LLM. The GoAI-CoT-Reviewer evaluation model is built on the Llama 3.1-8B-Instruct model (Dubey et al. 2024) and fine-tuned through LoRA. Training is conducted using NVIDIA A100 GPUs, with a LoRA rank set to 8.
Evaluation Method
As the only trainable component within the GoAI system, we first conducted an experimental evaluation of the correlation between the GoAI-CoT-Reviewer and human judgments, thereby establishing the reliability and practical utility of the GoAI framework as a whole and laying the groundwork for subsequent experiments.
To evaluate the overall effectiveness of the GoAI system, we recruited 9 upper-level undergraduate and early-stage graduate students majoring in AI, all possessing foundational domain knowledge, and assigned them to one of three groups in a between-subjects randomized design. Each group employed a distinct method to analyze research trends and generate ideas: (G1) web search combined with manual paper reading, (G2) interaction with an LLM-based tutoring assistant, and (G3) use GoAI Idea Studio. Participants were instructed to complete two tasks within a two-hour time frame: Frontier Mapping, which involved drafting a one-page summary including key baselines, extensions, contrasts, datasets/tools, and a list of prerequisite knowledge; and Creative Ideation, which required proposing one to two project ideas related to the assigned topic. Participants using LLM and GoAI were asked to complete the System Usability Scale (SUS), while all participants completed the NASA-TLX to assess task load. Additionally, we recorded the scores of the ideas generated by all participants.
Main Results
Human-Model Agreements of Evaluation
Table 2 presents the Pearson correlation coefficients between our evaluation model and human assessments. The results indicate that our model demonstrates a strong positive correlation with human reviewers, validating its effectiveness as a source of feedback in the idea generation process. Compared to models trained on raw data or multi-turn dialogue datasets, those incorporating a structured Chain-of-Thought (CoT) reasoning process exhibit significantly improved alignment with human evaluations, a finding that is consistent with the conclusions drawn in (Xu et al. 2025).
The Effectiveness of GoAI
Table 1 presents the performance of student groups employing different methods, highlighting the effectiveness of GoAI. In terms of system usability, the scores for GoAI and the LLM-based assistant are comparable, indicating that despite the greater complexity of the GoAI system, it does not impose additional operational burdens or steep learning curves on users. Regarding task-related cognitive load, students using GoAI reported lower levels of mental and time pressure, suggesting that the GoAI approach effectively supports AI students in synthesizing research trends, identifying prerequisite knowledge, and generating novel research ideas. Notably, students utilizing GoAI also outperformed their peers in the novelty of the ideas produced, demonstrating its capacity to enhance creative ideation in academic contexts.
Conclusion
In this paper, we present Graph of AI Ideas (GoAI), a novel framework that integrates LLMs with knowledge graphs to support AI students in systematically understanding research trends, constructing prerequisite-aligned learning pathways, and generating innovative research ideas. By organizing ideas, methodologies, and conceptual elements from existing literature into a structured graph representation, GoAI effectively captures the progressive relationships among research works. Performing path search within this graph enables the LLM to better comprehend and analyze ongoing advancements, uncover interconnections among studies, and thereby generate more coherent developmental trajectories and tailored learning paths. To evaluate and iteratively refine the automatically generated ideas, we introduce GoAI-CoT-Reviewer, a structured reasoning-based evaluation model that assesses the novelty of research ideas through a process closely aligned with human peer review, yielding feedback with high agreement to expert judgment. Experimental results demonstrate that GoAI offers strong usability, significantly reduces students’ cognitive load in synthesizing research landscapes and learning requirements, and effectively fosters the generation of higher-quality and more creative research ideas.